OpenStudy (anonymous):

The line with slope -2 passes through the point (-3,-3) . Find the equation of the line in standard form: Ax+By=C?

OpenStudy (dumbcow):

-3=(-2)(-3) +b b = -9 y = -2x -9 2x+y=-9

OpenStudy (yuki):

dumbcow used the formula for "linear equations" y=mx+b where m is the slope and b is the y-intercept. if you remember this all you have to do is to plug in all the x and y to solve. since y = mx+b is called the "slope intercept form" and Ax + By = C is called the "standard form" you have to fix it back by manipulating the equation.

OpenStudy (dumbcow):

thanks for clarifying